A coin lies at the bottom of a tank containing water to a depth of 130cm. If the refractive index is 1.3. Calculate the apparent depth of the coin.
1
Expert's answer
2018-11-05T15:53:09-0500

Real depth is H. Apparent depth is h.

n=H/h

1.3=130/h

h=130/1.3=100 cm.